OpenStudy (anonymous):

Trigonometry Hellppp please: [Given that sin θ= -4/5 with θ in quadrant IV, find cos 2θ.] --please show me how you got your answer, thank you :D

myininaya (myininaya):

\[\sin(\theta)=\frac{opp}{hyp} \text{ , right? }\]

OpenStudy (anonymous):

yes it does

myininaya (myininaya):

 So can you label this triangle with the info given \[\sin(\theta)=\frac{-4}{5}\]

myininaya (myininaya):

I put that triangle in the 4th quadrant because that is where it told me to put it

myininaya (myininaya):

You can use my picture just click that pencil and label away

OpenStudy (anonymous):

ohhhhhhh it is that simple....hah i think i over analyzed it...thanks will do